# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v2

@@ Appendix A Limitations: Computational costs
The improved extrapolation accuracy offered by the proposed TI-DeepONet and TI(L)-DeepONet approaches comes at the cost of increased training and inference times. This additional computational overhead stems from the repeated forward and backward passes through the time integrator. However, the convergence of TI-based networks is significantly faster, indicating more efficient optimization of network parameters compared to baseline methods. Table A1 presents a detailed comparison of the training and inference costs for each method across all benchmark problems. Moreover, incorporating adaptive, learnable slope coefficients facilitates convergence to the true solution even faster due to the solution-specific local adaptivity within the time integrator. Training and inference for the 1D Burgers’, 1D KdV, and 1D KS cases were conducted on a single NVIDIA A100 GPU equipped with 40GB memory (Intel Xeon Gold Cascade Lake). The 2D Burgers’ case was performed on a NVIDIA A100 GPU with 80GB memory (Intel Xeon Gold Icy Lake).
@@
Table A1: Computational costs of training and inference for all examples across different frameworks.
@@
During training, the computational complexity varies across methods. The branch network requires $\mathcal{O}(bs)$ forward passes per iteration, where $bs$ represents the batch size. The trunk network evaluation complexity depends on the strategy employed:
@@
  • Autoregressive methods (autoregressive DeepONet, TI-DeepONet, and TI(L)-DeepONet): $\mathcal{O}(n_{s})$ forward passes because of the absence of temporal dimension.
For TI(L)-DeepONet, an additional $\mathcal{O}(bs)$ forward passes are required per iteration to evaluate the auxiliary deep neural network that learns the intermediate RK4 slopes. As a result, TI(L)-DeepONet incurs the highest per-iteration cost, leading to the lowest iteration speed observed in Table A1. Although TI-DeepONet and autoregressive DeepONet evaluate a similar number of forward passes, TI-DeepONet additionally invokes the time integrator, introducing extra forward computations. Consequently, it is more expensive to train than autoregressive DeepONet but still cheaper than TI(L)-DeepONet. The cost of the full rollout method scales directly with the spatiotemporal resolution. While it avoids recursion, the simultaneous prediction over all space-time points makes it more expensive than autoregressive methods, which predicts for only spatial points at each step. The trends observed in Table A1 align with this analysis. Note that in the 2D Burgers’ case, the relatively higher training cost incurred by the autoregressive DeepONet is due to the use of additional hidden layers in the trunk network.
@@
Appendix B Statistical Variability
To evaluate the performance of the model, we compute the relative $L_{2}$ error of the extrapolation predictions, and report the mean and standard deviation of this metric based on five independent training trials in Table A2.
@@ Appendix B Statistical Variability
Across five independent trial runs, TI(L)-DeepONet consistently achieves the lowest extrapolation error across all benchmark problems. In addition to superior accuracy, it also exhibits reduced variance in error estimates, indicating enhanced robustness and reliability. As expected, both the full rollout and autoregressive DeepONet models suffer from substantially higher errors, particularly outside the training temporal domain. Similarly, the FNO variants, both full rollout and autoregressive incur higher errors. FNO FR generally exhibits higher errors throughout the entire temporal domain and yields particularly unstable solutions due to its inability to capture all global frequencies of the system. Notably, this limitation also manifests as higher variance in the predictions. For FNO AR, spectral instability causes error spikes in the extrapolation domain, leading to higher solution variance - consistent with our observations highlighted in Sec. 3. Overall, the TI-based frameworks not only enhance extrapolation accuracy but also offer greater robustness across random initializations, underscoring their effectiveness for modeling time-dependent PDEs.
Appendix C Additional details
C.1 One-dimensional Burgers’ Equation
Data generation:
Data preparation:
The training and testing datasets are constructed differently depending on the learning strategy employed. In the full rollout setting, the goal is to train a DeepONet/FNO to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps an initial condition $s(x)$ to the full spatiotemporal solution $u(x,t)$. For this purpose, we consider a set of initial conditions, spatiotemporal query locations, and the corresponding solution fields. The data is partitioned into training and testing sets using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2000$ and $N_{\text{test}}=500$ samples. Importantly, we aim to assess the performance of various frameworks in learning the full spatiotemporal field over the entire temporal domain when given access to solutions at limited timesteps. To this end, we restrict training to only half of the temporal domain, i.e., $t_{\text{train}}\in[0,0.5]$. Consequently, the grid for the trunk network is formed by taking a meshgrid over spatiotemporal coordinates $(x,t)$, where $x\in[0,1]$ and $t\in[0,0.5]$ during training, while $t\in[0,1]$ during testing. This same meshgrid is also employed for computing the 2D Fourier Transform in FNO full rollout.
For training the autoregressive DeepONet, autoregressive FNO, TI-DeepONet, and TI(L)-DeepONet architectures, we adopt a different strategy for preparing the training and testing datasets. In autoregressive methods, future states of the system are recursively predicted using the model’s own previous outputs as inputs at each timestep. This approach inherently leads to error accumulation, as prediction errors at one timestep propagate and potentially amplify at subsequent timesteps. Understanding and mitigating this issue is a key motivation behind the TI-DeepONet architecture. To this end, the training dataset is created as follows: the input to the branch network consists of stacked solution states from $t=0$ to $t=0.5$, i.e., $u^{0}(x),u^{1}(x),\ldots,u^{50}(x)]$. Similarly, the output ground truth consists of stacked solution states from $t=0.01$ to $t=0.51$, i.e., $u^{1}(x),u^{2}(x),\ldots,u^{51}(x)]$. This arrangement enables the solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ to learn the mapping $u^{i}(x)\rightarrow u^{i+1}(x)$, where $u^{i}(x)$ and $u^{i+1}(x)$ represent the solution states at the $i^{\text{th}}$ and $(i+1)^{\text{th}}$ timesteps, respectively. The trunk network processes only the spatial coordinates $x\in[0,1]$. For autoregressive FNO, the inputs and outputs are distinct $u^{i}(x)$ and $u^{i+1}(x)$ solution pairs, and the Fourier transform is one-dimensional and is applied only in the $x$-dimension.
Training architecture:
Tables A3 and A4 present the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet architecture, we additionally employ an auxiliary feedforward network designed to predict the four RK4 slope coefficients. This network consists of two hidden layers with 32 neurons each, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate scheduling strategy was employed for the auxiliary network predicting the adaptive RK4 slope coefficients. Notably for FNO AR, the learning rate schedule also employed exponential decay but by a factor of 0.96 every 2000 steps. Both training and test losses were monitored at each epoch. The final model corresponds to the set of parameters that yielded the lowest test loss over the entire training process. Among all tested architectures, TI(L)-DeepONet exhibited the fastest convergence, attributed to its capacity to adapt to the local dynamics of the solution. Later in this section, we examine the evolution of the learned RK4 slope coefficients across epochs and compare them to the reference coefficients: $\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}$.
@@ C.1 One-dimensional Burgers’ Equation
Table A3: 1D Burgers’ equation: Neural network architecture and training settings across different DeepONet-based methods.
@@
As previously discussed, the learnable slope coefficients $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]$ are predicted by the additional NN, which takes the current solution state $u^{i}$ as input, i.e., $\boldsymbol{\alpha}=\text{NN}_{\theta}(u^{i})$. Analyzing the evolution of these solution-specific coefficients over training epochs provides insight into how the network adapts the time integrator to local dynamics, in contrast to using a fixed set of RK4 coefficients. This comparison is illustrated in Fig. A1. The dashed blue line in each plot indicates the reference slope coefficients typically used in the RK4 method, derived from a Taylor series expansion. These coefficients result in a fourth-order accurate scheme and are generally expected to provide high accuracy for traditional time integrators. However, in the context of the operator learning framework, these coefficients also adapt to the approximation error of the neural operator. Making these slope coefficients learnable allows the model to better compensate for such errors, thereby improving the overall accuracy of the framework.
C.2 One-dimensional Korteweg-de Vries (KdV) Equation
Data generation:
$u_{i}(x,0)=2k_{i}^{2}sech^{2}\left(k_{i}\left(\left((x+\frac{P}{2}-Pd_{i}\right)\%P-\frac{P}{2}\right)\right)^{2}$ (10)
where sechsech is hyperbolic secant, $P$ is the period in space, % is the modulo operator, $i=\{1,2\}$, $k\in[0.5,1]$ and $d\in[0,1]$ are the coefficients that determine the height and location of the soliton peak, respectively. The dataset is curated by generating $N=1000$ distinct initial conditions and their temporal evolution is modeled using the midpoint method. The 1D spatial domain, $\Omega=[0,10]$, is discretized into 100 uniformly spaced grid points, resulting in a spatial resolution of $\Delta x=0.1$. For the temporal domain, we consider a uniform grid of 201 points spanning the time domain $t\in[0,5]$, which yields a $\Delta t=0.025$.
Data preparation:
Training architecture:
Tables A5 and A6 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet framework, we employ a secondary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 and 48 neurons, respectively, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure that the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate schedule was used for training the auxiliary network tasked with learning the RK4 slope coefficients. The same learning rate schedule with identical hyperparameters was employed for FNO FR training. For FNO AR, an exponential decay learning rate schedule was also used, but with an initial learning rate of $10^{-3}$ that decayed by a factor of 0.96 every 2000 steps.
Table A5: 1D KdV equation: Neural network architecture and training settings across different DeepONet methods.
@@ C.2 One-dimensional Korteweg-de Vries (KdV) Equation
C.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Data generation:
We consider the 1D Kuramoto-Sivashinsky (KS) equation on a spatial domain $x\in[0,L]$ with $L=6\pi$, equipped with periodic boundary conditions. The initial conditions $u(x,0)=s(x)$ are sampled from mean-zero Gaussian random fields (GRFs) with periodic boundary conditions. Specifically, we generate GRFs with covariance operator $\mathcal{C}=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$, where $\gamma=1$, $\tau=2$, and $\sigma=1$. The spatial domain is discretized using $N_{x}=128$ grid points, and the temporal domain $t\in[0,30]$ is discretized with $N_{t}=300$ timesteps. The PDE is solved using the exponential time-differencing fourth-order Runge-Kutta (ETDRK4) method with a timestep of $\Delta t=0.1$. We generate a total of 3000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(x,t)$. The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2400$ training samples and $N_{\text{test}}=600$ testing samples. The periodic boundary conditions ensure that the solution remains in the space of mean-zero periodic functions throughout the evolution, preserving the chaotic dynamics characteristic of the KS equation.
Data preparation:
Training architecture:
Tables A7 and A8 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the full rollout DeepONet, both spatial and temporal coordinates $(x,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 2 to 42. For the autoregressive and TI-based variants, only the spatial coordinate $x$ is augmented with Fourier features $\sin(\pi kx),\cos(\pi kx)]$, expanding the input dimension from 1 to 21. For the TI(L)-DeepONet framework, we employ an auxiliary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 neurons each, followed by an output layer with four neurons. The hidden layers utilize the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. All DeepONet variants (DON FR, DON AR, TI-DON, and TI(L)-DON) are trained using the AdamW optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.95 every 5000 epochs, and a weight decay of $10^{-4}$. For TI(L)-DeepONet specifically, the auxiliary RK4 network is trained separately using the Adam optimizer with an initial learning rate of $2\times 10^{-3}$ and the same exponential decay schedule. The FNO variants employ the Adam optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.96 every 2000 steps for both FNO FR and FNO AR configurations.
Table A7: 1D KS equation: Neural network architecture and training settings across different DeepONet methods.

### v3

@@ 4 Computational Costs
The improved extrapolation accuracy offered by the proposed TI-DeepONet and TI(L)-DeepONet approaches comes at the cost of increased training and inference times. This additional computational overhead stems from the repeated forward and backward passes through the time integrator. However, the convergence of TI-based networks is significantly faster, indicating more efficient optimization of network parameters compared to baseline methods. Table 2 presents a detailed comparison of the training and inference costs for each method across all benchmark problems. Moreover, incorporating adaptive, learnable slope coefficients facilitates convergence to the true solution even faster due to the solution-specific local adaptivity within the time integrator. Training and inference for the 1D Burgers’, 1D KdV, and 1D KS cases were conducted on a single NVIDIA A100 GPU equipped with 40GB memory (Intel Xeon Gold Cascade Lake). The 2D Burgers’, 2D Rotation Advection-Diffusion, and 3D Heat Conduction cases were performed on a NVIDIA A100 GPU with 80GB memory (Intel Xeon Gold Icy Lake).
@@
Table 2: Computational costs of training and inference for all examples across different frameworks. Note that training speed is reported as iterations per second, where a lower value indicates slower training and higher computational cost.
@@
Rotation Advection Diffusion (2D) TI(L)-DON (Ours) 32 6.27 800 7.54
TI-DON AB (Ours) 6.89 4.89
DON Full rollout 55.43 2.48
DON Autoregressive 32.55 5.37
FNO Full rollout 1.67 21.76
FNO Autoregressive 65.19 15.84
Heat Conduction (3D) TI(L)-DON (Ours) 64 10.64 1000 14.39
TI-DON AB (Ours) 12.83 11.35
DON Full rollout 2.99 2.62
DON Autoregressive 55.44 11.47
@@
For the sake of comparison of the theoretical computational complexity for the TI-based variants, since the backbone NO is DeepONet, we present the theoretical comparison with respect to DON FR and DON AR. We note that a similar analysis can be performed agnostic of the specific NO architecture in a more general setting. During training, the computational complexity varies across methods. The branch network requires $\mathcal{O}(bs)$ forward passes per iteration, where $bs$ represents the batch size. The trunk network evaluation complexity depends on the strategy employed:
@@
  • Autoregressive methods (autoregressive DeepONet, TI-DeepONet, and TI(L)-DeepONet): $\mathcal{O}(n_{s})$ forward passes due to the absence of the temporal dimension.
For TI(L)-DeepONet, an additional $\mathcal{O}(bs)$ forward passes are required per iteration to evaluate the auxiliary deep neural network that learns the intermediate RK4 weighting coefficients. As a result, TI(L)-DeepONet incurs the highest per-iteration cost, leading to the lowest iteration speed observed in Table 2. Although TI-DeepONet and autoregressive DeepONet evaluate a similar number of forward passes, TI-DeepONet additionally invokes the time integrator, introducing extra forward computations. Consequently, it is more expensive to train than autoregressive DeepONet but still cheaper than TI(L)-DeepONet. The cost of the full rollout method scales directly with the spatiotemporal resolution. While it avoids recursion, the simultaneous prediction over all space-time points makes it more expensive than autoregressive methods, which predict for only spatial points at each step. The trends observed in Table 2 align with this analysis. Note that in the 2D Burgers’ and 2D Rotation Advection-Diffusion cases, the relatively higher training cost incurred by the autoregressive DeepONet is due to the use of additional hidden layers in the trunk network.
@@
5 Choice of Inference Timestep
In our primary evaluations, TI-DeepONet was trained and tested using the same temporal resolution, i.e., the training and inference timesteps ($\Delta t$) were matched. However, the architecture’s compatibility with higher-order numerical integrators suggests that, in principle, inference may be conducted using timesteps different from those used during training. As long as the chosen $\Delta t$ satisfies the Courant-Friedrichs-Lewy (CFL) stability condition, the resulting time integration should remain stable. This flexibility stems from the fact that TI-DeepONet mirrors traditional numerical solvers: both employ explicit time-stepping schemes, with TI-DeepONet approximating the right-hand side of the PDE via a neural operator rather than finite-difference or finite-element discretizations of spatial derivatives.
To investigate this capability, we performed preliminary experiments on the 1D Burgers’ equation, evaluating TI-DeepONet at three inference timesteps, $\Delta t\in\{0.1,0.01,0.001\}$, while training was performed at $\Delta t=0.01$. Notably, TI-DeepONet preserved comparable $L_{2}$ error even when evaluated with a significantly coarser timestep ($\Delta t=0.1$) than that used during training. This observation is further corroborated by the 2D Burgers’ case, as shown in Figure 15, where both TI-DON and TI(L)-DON maintain consistent performance across varying inference timesteps. This finding suggests the potential for developing more efficient surrogates that accelerate inference without additional training costs.
6 Conclusions
Accurate long-term extrapolation of dynamical systems remains a fundamental challenge for neural operators, particularly due to error accumulation in autoregressive predictions. To address this, we introduce TI-DeepONet and its adaptive variant TI(L)-DeepONet, which integrates classical numerical time integration with an operator learning framework (DeepONet). By reformulating the learning objective to approximate instantaneous time derivatives—subsequently integrated via high-order numerical schemes, our frameworks preserve the causal structure of dynamical systems while enhancing stability and accuracy in long-term forecasting. Furthermore, the operator learning paradigm enables continuous spatiotemporal solutions for parametrized initial conditions, thereby enabling the construction of an efficient and reliable emulator for complex systems. Extensive evaluations across six canonical PDEs (1D/2D Burgers’, 1D KdV, 1D KS, 2D Rotation Advection-Diffusion, and 3D Heat Conduction equations) demonstrate the effectiveness of our approach. For the majority of benchmark problems, particularly diffusion-dominated and chaotic systems, the TI-based variants demonstrate superior performance over baseline methods. TI-DeepONet reduces relative $L_{2}$ extrapolation errors by 96.3% compared to autoregressive DeepONet and 83.6% compared to full rollout methods, while maintaining stable predictions for temporal domains extending up to twice the training interval. The learnable coefficients in TI(L)-DeepONet further refine accuracy by adaptively weighting intermediate integration slopes, enabling solution-specific adjustments that outperform static numerical schemes in most of our examples. However, for the rotation advection-diffusion problem involving rigid body rotational dynamics, FNO AR outperforms the TI-based variants, suggesting that the choice of neural operator architecture may be as important as the time integration strategy for certain dynamical systems. This motivates a broader perspective: while we proposed TI-DeepONet with DeepONet as the neural operator in this work, the framework can be viewed more generally as TI-NO (Time-Integrator embedded Neural Operator), where both the time integration scheme and the neural operator architecture can be independently selected based on the problem characteristics. By embedding temporal causality into the learning process and decoupling training-time integrators (e.g., Runge-Kutta) from inference-time integrators (e.g., Runge-Kutta or Adams-Bashforth), our methodology bridges neural operator flexibility with the robustness of numerical analysis. Among the proposed variants, TI(L)-DeepONet achieves the highest accuracy, albeit with a higher per-iteration computational cost due to its learnable integration coefficients (see SI Section 4). However, this increase is offset by its significantly faster convergence to lower training loss, resulting in no net increase in total compute time. This hybrid paradigm offers a reliable pathway for modeling high-dimensional, time-dependent systems, where stability and generalizability are critical, though problem characteristics should guide the selection of the most appropriate method and neural operator architecture.
7 Limitations and Future Work
While our proposed frameworks demonstrate superior extrapolation accuracy compared to existing neural operator training strategies, they incur higher computational costs during both training and inference, as detailed in Section 4. Although these offline training costs can be justified by the consistent reliability of predictions across a wide range of parametric initial conditions, improving computational efficiency remains a key priority. To this end, we plan to leverage multi-GPU training and investigate methods for more efficient forward and backward propagation through the ODE solver. Additionally, we aim to extend the framework’s scalability to more complex 3D and multiphysics systems and to incorporate adaptive time-stepping strategies for handling stiff dynamical systems. Furthermore, as evidenced by the rotation advection-diffusion example, the TI-based variants exhibit limitations in capturing rigid body rotational dynamics; improving performance for such problems through alternative neural operator architectures (e.g., TI-FNO) or finer temporal resolution strategies remains an important direction for future investigation.
Acknowledgments
Appendix A Statistical Variability
@@
To evaluate the performance of the model, we compute the relative $L_{2}$ error of the extrapolation predictions, and report the mean and standard deviation of this metric based on five independent training trials in Table A1. Across five independent trial runs, TI(L)-DeepONet consistently achieves the lowest extrapolation error across all benchmark problems. In addition to superior accuracy, it also exhibits reduced variance in error estimates, indicating enhanced robustness and reliability. As expected, both the full rollout and autoregressive DeepONet models suffer from substantially higher errors, particularly outside the training temporal domain. Similarly, the FNO variants, both full rollout and autoregressive incur higher errors. FNO FR generally exhibits higher errors throughout the entire temporal domain and yields particularly unstable solutions due to its inability to capture all global frequencies of the system. Notably, this limitation also manifests as higher variance in the predictions. For FNO AR, spectral instability causes error spikes in the extrapolation domain, leading to higher solution variance - consistent with our observations highlighted in Sec. 3. Overall, the TI-based frameworks not only enhance extrapolation accuracy but also offer greater robustness across random initializations, underscoring their effectiveness for modeling time-dependent PDEs.
Appendix B Additional details
B.1 One-dimensional Burgers’ Equation
@@
Table A2: 1D Burgers’ Equation: Neural network architecture and training settings across different DeepONet-based methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.
@@
As previously discussed, in the TI(L)-DeepONet architecture, the learnable slope coefficients $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]$ are predicted by the additional NN, which takes the current solution state $u^{i}$ as input, i.e., $\boldsymbol{\alpha}=\text{NN}_{\theta}(u^{i})$. Analyzing the evolution of these solution-specific coefficients over training epochs provides insight into how the network adapts the time integrator to local dynamics, in contrast to using a fixed set of RK4 coefficients. This comparison is illustrated in Fig. A1. The dashed blue line in each plot indicates the reference slope coefficients typically used in the RK4 method, derived from a Taylor series expansion. These coefficients result in a fourth-order accurate scheme and are generally expected to provide high accuracy for traditional time integrators. However, in the context of the operator learning framework, these coefficients also adapt to the approximation error of the neural operator. Making these slope coefficients learnable allows the model to better compensate for such errors, thereby improving the overall accuracy of the framework.
B.2 One-dimensional Korteweg-de Vries (KdV) Equation
Table A4: 1D KdV Equation: Neural network architecture and training settings across different DeepONet methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.
@@ B.2 One-dimensional Korteweg-de Vries (KdV) Equation
B.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Table A6: 1D KS Equation: Neural network architecture and training settings across different DeepONet methods. Note that each epoch corresponds to a single gradient descent update over a randomly sampled batch, equivalent to one stochastic gradient descent iteration.